OpenStudy (dtan5457):

What is the empirical formula of an ion whose composition by mass is 57.14% sulfur and 42.86% oxygen?

OpenStudy (dtan5457):

All the ions have a charge of -2.

OpenStudy (dtan5457):

For problems like this I just get the moles and assume the total is 100g.. 57.14/32 and 42.86/16 however I get 1.79 and 2.677...not sure what to do

OpenStudy (dtan5457):

@Zale101

OpenStudy (dtan5457):

@aaronq

OpenStudy (zale101):

Assume the total of grams of compound is 100g. So, there would be 57.14g of sulfur and 42.86g of oxygen.

OpenStudy (dtan5457):

yeah

OpenStudy (zale101):

Convert to moles and those moles will be in the element's subscript.

OpenStudy (dtan5457):

I get 1.79 and 2.677

OpenStudy (zale101):

Good, now divide the subscript all by the lowest mole.

OpenStudy (zale101):

1.79 is for sulfur, right?

OpenStudy (dtan5457):

yes

OpenStudy (zale101):

1.79 mols*

OpenStudy (dtan5457):

like 1.5

OpenStudy (dtan5457):

for f

OpenStudy (dtan5457):

I mean by 2. So I get a whole number

OpenStudy (zale101):

If you get a number such as 1.4, then you'd have to multiply it by 2 to all the subscript (sulfur and oxygen) to get a whole number. If it comes out to be 2.9 that's fine because that is closer to 3 (a whole number).

OpenStudy (dtan5457):

So I'm looking at S2F3^-2

OpenStudy (zale101):

\(\large S_2F_3\)

OpenStudy (zale101):

^^ That would be the empirical formula

OpenStudy (dtan5457):

the -2 is just the charge since it's a ion

OpenStudy (dtan5457):

according to the question

OpenStudy (zale101):

S is sulfur and and F is fluorine and both are nonmetals, so they are covalently bonded together (both of them have negative charges and that's why they share electrons).

OpenStudy (zale101):

When dealing with writing chemical formulas for covalent compounds, make sure you don't focus on the charges because there all negatives and there's no opposite signs that makes them want to balance as + & -

OpenStudy (zale101):

balance out*

OpenStudy (zale101):

makes sense?

OpenStudy (dtan5457):

Wouldn't the charge still have to be 0?

OpenStudy (zale101):

but all charged elements are ions, including negatives but if a negative element is boned with a negative element , there charges will share and they wont play as "ions" no more.

OpenStudy (zale101):

The charge will have to be zero ONLY if there's an ionic or acids compounds.

OpenStudy (zale101):

Like NaCl, are ionic compounds because there's 1 cation (+) and 1 anion (-) Na (sodium) has 1+ charge Cl (chlorine) has 1- charge 1-1=0

OpenStudy (zale101):

but we are dealing with S2F3 and those are covalent and not ionic

OpenStudy (dtan5457):

Yeah, I see, like in my book when I was learning covalent we had to find the oxidation numbers, but the charge was never 0.

OpenStudy (zale101):

I see what you're getting at. Well, with oxidation states, i dont think we usually include them in this question.

OpenStudy (dtan5457):

I know, it's just that all 4 answer choices have a charge of -2 lol

OpenStudy (zale101):

OOPs it's oxygen not Florine geez, sorry about the mistake

OpenStudy (zale101):

I'm blind lol

OpenStudy (dtan5457):

Oh my..lol. I was using florine too..I was kind of wondering why the charge would ever be -2..

OpenStudy (zale101):

\(\LARGE S_{1.79} ~:~O_{2.68}\) \(\LARGE S_{\frac{1.79}{1.79}} ~:~O_{\frac{2.68}{1.79}}\) times by 2 to the subscripts to get whole numbers \(\LARGE S_{2} ~:~O_{3}\)

OpenStudy (zale101):

Yes, then you were right about what you said regarding oxidation states...

OpenStudy (zale101):

Why Fluorine Zale???? lol

OpenStudy (dtan5457):

Which actually leads to another question...o3=-6 charge?

OpenStudy (dtan5457):

O is diatomic and written as o2...is that -4 or -2?

OpenStudy (zale101):

there's 2 sulfure, but the charge we don't know (x) and we know the oxygen is 2- and it has3 mols. Also, in oxidation state, the first element has a positive charge. 2 S x (3+)- 3 O (2-)=0 +6-6=0 \(O_3\) is -6 which means that \(S_2\) must be +6

OpenStudy (zale101):

2 S *(3+)- 3 O *(2-)=0

OpenStudy (dtan5457):

Why is the charge -2 then?

OpenStudy (zale101):

I dont have a good explanation for this. Maybe @aaronq can helps us out.

OpenStudy (dtan5457):

@Abhisar Apparently S2O3^-2 is thiosulfate. regular s2o3 is disulfur trioxide. so...i guess if they want an ion, the ^-2 will work